# Robust Amortized Bayesian Inference with Self-Consistency Losses on Unlabeled Data

### v4

@@ Abstract
Amortized Bayesian inference (ABI) with neural networks can solve probabilistic inverse problems orders of magnitude faster than classical methods. However, ABI is not yet sufficiently robust for widespread and safe application. When performing inference on observations outside the scope of the simulated training data, posterior approximations are likely to become highly biased, which cannot be corrected by additional simulations due to the bad pre-asymptotic behavior of current neural posterior estimators. In this paper, we propose a semi-supervised approach that enables training not only on labeled simulated data generated from the model, but also on unlabeled data originating from any source, including real data. To achieve this, we leverage Bayesian self-consistency properties that can be transformed into strictly proper losses that do not require knowledge of ground-truth parameters. We test our approach on several real-world case studies, including applications to high-dimensional time-series and image data. Our results show that semi-supervised learning with unlabeled data drastically improves the robustness of ABI in the out-of-simulation regime. Notably, inference remains accurate even when evaluated on observations far away from the labeled and unlabeled data seen during training.
@@ 1 Introduction
Theory-driven computational models (mechanistic models) are highly influential across numerous branches of science [29]. The utility of computational models largely stems from their ability to fit real data $x$ and extract information about hidden parameters $\theta$. Bayesian methods have been instrumental for this task, providing a principled framework for uncertainty quantification and inference [18]. However, gold-standard Bayesian methods, such as Gibbs or Hamiltonian Monte Carlo samplers [3], remain notoriously slow. Moreover, these methods are rarely feasible for fitting complex models [5] or even simpler models in big data settings with many thousands of data points in a single dataset [2], or when thousands of independent datasets require repeated model re-fits [42].
@@
In recent years, deep learning methods have helped address some of these efficiency challenges [4]. In particular, amortized Bayesian inference [ABI; 19, 30, 22, 37, 20, 9, 46] has received considerable attention for its potential to automate Bayesian workflows by training generative neural networks on model simulations, subsequently enabling near-instant downstream inference on real data. However, due to the reliance on pre-trained neural networks, ABI methods can become unreliable when applied to data that is unseen or sparsely encountered during training. In particular, posterior samples from amortized methods may deviate significantly from samples obtained with gold-standard MCMC samplers when there is a mismatch between the simulated training data and the real data [43, 38, 40, 20, 16]. This lack of robustness limits the widespread and safe applicability of ABI methods.
@@
In this work, we propose a new robust semi-supervised approach to ABI. The supervised part learns from a “labeled” set of parameters and corresponding synthetic (simulated) observations, $\{\theta,x\}$, while the unsupervised part leverages an “unlabeled” data set of real observations $\{x^{*}\}$ without parameters. In contrast to other methods aiming to enhance the robustness of ABI, our approach does not require ground-truth parameters $\theta^{*}$ [44], post hoc corrections [43, 40], or specific adversarial defenses [20], nor does it entail a loss of amortization [43, 25] or generalized Bayesian inference [17, 35].
To achieve robust inference, we expand on previous work on self-consistency losses [39, 26] and demonstrate notable robustness gains even for as few as four real-world observations. We provide theoretical proofs for the strict properness of our semi-supervised approach based on self-consistency. We also show that self-consistency losses can be added to any standard simulation-based objective without introducing trade-offs. Empirical results on a variety of tasks including high-dimensional time-series and image data demonstrate that our approach retains ABI’s characteristic speed while achieving remarkable robustness: posterior estimates remain accurate and well-calibrated even for observations different from both the labeled and unlabeled training data. By unifying theoretical guarantees with practical performance, our method represents a significant step toward safe and reliable ABI in the presence of simulation gaps.
@@ 2.1 Bayesian self-consistency
Self-consistency leverages a simple symmetry in Bayes’ rule to enforce more accurate posterior estimation even in regions with sparse data [39, 26]. Crucially, it incorporates likelihood (when available) or a surrogate likelihood during training, thereby providing the networks with additional information beyond the standard simulation-based loss typically employed in ABI (see below).
Following [39], we will focus on the marginal likelihood based on neural posterior or likelihood approximation. Under exact inference, the marginal likelihood is independent of the parameters $\theta$. That is, the Bayesian self-consistency ratio of likelihood-prior product and posterior is constant across any set of parameter values $\theta^{(1)},\dots,\theta^{(L)}$,
@@
However, replacing $p(\theta\mid x)$ with a neural estimator $q(\theta\mid x)$ (likewise for the likelihood) leads to undesired variance in the marginal likelihood estimates across different parameter values on the right-hand-side [39]. Since this variance is a proxy for approximation error, we can directly minimize it via backpropagation along with any other ABI loss to provide further training signal and reduce errors guided by density information. Our proposed semi-supervised formulation builds on these advantageous properties.
@@ 2.2 Semi-supervised amortized Bayesian inference
The formulation in Eq. (1) is straightforward, but practically never used in traditional sampling-based methods (e.g., MCMC) because they do not provide a closed-form for the approximate posterior density $q(\theta\mid x)$. In contrast, we can readily evaluate $q(\theta\mid x)$ in ABI when using a neural density estimator that allows efficient density computation (e.g., normalizing flows, [28]). Thus, we can formulate a family of semi-supervised losses of the form:
@@
where $S$ is a strictly proper score [21] and $C$ is a self-consistency score [39]. The neural networks to be optimized are a generative model $q$ and (potentially) a summary network $h$ extracting lower dimensional sufficient statistics from the data. We will call the first loss component, $\mathbb{E}_{(\theta,x)\sim p(\theta,x)}\left[S(q(\theta\mid h(x)),\theta)\right]$, the (standard) simulation-based loss, as it forms the basis for standard ABI approaches using simulation-based learning. E.g., this is the maximum likelihood loss for normalizing flows [28, 36] or a vector-field loss for flow matching [33, 32]. We will refer to the second loss component as the (Bayesian) self-consistency loss.
@@
$(q^{*},h^{*})=\operatorname*{argmin}_{q,h}\frac{1}{N}\sum_{n=1}^{N}\left[S(q(\theta_{n}\mid h(x_{n})),\theta_{n})\right]+\lambda\cdot\frac{1}{M}\sum_{m=1}^{M}\left[C\left(\frac{p(x^{*}_{m}\mid\theta)\,p(\theta)}{q(\theta\mid h(x^{*}_{m}))}\right)\right].$ (3)
Asymptotically for $N\rightarrow\infty$, that is, for infinite training data generated from the simulator $p(\theta,x)$, a universal density estimator [7] minimizing a strictly proper simulation-based loss [21] is sufficient to ensure perfect posterior approximation for any data. By this, we mean that the posterior approximation becomes identical to the posterior we would obtain if we could analytically compute $p(\theta\mid x)=p(x\mid\theta)p(\theta)/p(x)$. This analytic posterior is sometimes also referred to as “true” or “correct” posterior. In practice, the posterior is rarely analytic, but we can still verify the accuracy of an approximation by comparing it with the results of a gold-standard approach (if available), such as a sufficiently long, converged MCMC run [34].
While neural posterior approximation is perfect asymptotically, its pre-asymptotic performance, that is, when training $q(\theta\mid h(x))$ only on a finite amount of simulated data, can become arbitrarily bad [16, 38]. For any data $x^{*}$ that is outside the data space implied by $p(\theta,x)$, for instance, when the model is misspecified, the posterior approximation $q(\theta\mid h(x^{*}))$ may be arbitrarily far away from the analytic posterior $p(\theta\mid x^{*})$. As a result, a simulation-based loss is insufficient to achieve robust ABI in practice. This is where the self-consistency loss comes in: As we will show, the latter greatly improves generalization to atypical data at inference time.
One particular choice for $C$ is the variance over parameters on the log scale of the Bayesian self-consistency ratio [39]:
@@
where $p_{C}(\theta)$ can be any proposal distribution over the parameter space, for example, the prior $p(\theta)$ or even the current approximate posterior $q_{t}(\theta\mid h(x^{*}))$ as given in a training iteration or snapshot $t$. Notably, the choice of $p_{C}(\theta)$ can influence training dynamics considerably, with the empirical consequences being difficult to anticipate [39]. In pratice, we approximate the variance $\text{Var}_{\theta\sim p_{C}(\theta)}$ by the empirical variance $\text{Var}_{l=1}^{L}$ computed over $L$ samples $\theta^{(l)}\sim p_{C}(\theta)$.
@@ 2.3 Self-consistency losses are strictly proper
Below, we discuss the strict properness of Bayesian self-consistency losses, which underline their widespread usefulness. To simplify the notation, we denote posterior approximators simply as $q(\theta\mid x)$ without considering architectural details such as the use of summary networks $h(x)$. All theoretical results and their proofs remain the same if $x$ is replaced by $h(x)$ as long as the summary network is expressive enough to learn sufficient statistics from $x$.
@@
The proofs of Propositions 1 and 2 are provided in Appendix A. The strict properness extends to semi-supervised losses of the form (2), which combine standard simulation-based losses with self-consistency losses.
@@
The proof of Proposition 3 follows immediately from the fact the sum of strictly proper losses is strictly proper. Importantly, since Proposition 3 holds independently of $p^{*}(x)$, it holds both in the case of a well-specified model, where $p^{*}(x)=p(x)$, and also in case of any model misspecification or domain shift where $p^{*}(x)\neq p(x)$. That is, there is no trade-off in the semi-supervised loss (2), since both loss components are both globally minimized for the same target.
Lastly, for completeness, we can also define strictly proper self-consistency losses for likelihood instead of posterior approximations.
@@
The proof of Proposition 4 proceeds in the same manner as for Proposition 1, just exchanging likelihood and posterior. Clearly, strict properness does not necessarily hold if both posterior and likelihood are unknown or approximate. This is because any pair of approximators $q(\theta\mid x)$ and $q(x\mid\theta)$ that satisfy $q(\theta\mid x)\propto q(x\mid\theta)\,p(\theta)$ minimize the self-consistency loss. For example, the choices $q(\theta\mid x)=p(\theta)$ and $q(x\mid\theta)\propto 1$ minimize the self-consistency loss, but may be arbitrarily far away from their actual target distributions $p(\theta\mid x)$ and $p(x\mid\theta)$, respectively.
In other words, if both likelihood and posterior are unknown, the self-consistency loss has to be coupled with another loss component, such as the maximum likelihood loss, to enable joint learning of both approximators $q(\theta\mid x)$ and $q(x\mid\theta)$ [39]. Nevertheless, the self-consistency loss still yields notable improvements: in our experiments, the semi-supervised loss (2) considerably enhanced the robustness of ABI even when both the posterior and likelihood are unknown.
@@ 3 Related work
The robustness of ABI and simulation-based inference methods more generally has been the focus of multiple recent studies [e.g., 15, 14, 16, 6, 43, 20, 25, 17, 40, 27, 44, 35, 38]. These efforts can be broadly classified into two categories: (a) analyzing or detecting simulation gaps and (b) mitigating the impact of simulation gaps on posterior estimates.
Since our work falls into the latter category, we briefly discuss methods aimed at increasing the robustness of fully amortized approaches. E.g., Gloeckler et al. [20] explore efficient regularization techniques that trade off some posterior accuracy to enhance the robustness of posterior estimators against adversarial attacks. Ward et al. [43] and Siahkoohi et al. [40] apply post hoc corrections based on real data, utilizing MCMC and the reverse Kullback-Leibler divergence, respectively.
Differently, Gao et al. [17] propose a departure from standard Bayesian inference by minimizing the expected distance between simulations and observed data, akin to generalized Bayesian inference with scoring rules [35]. Perhaps the closest work in spirit to ours is Wehenkel et al. [44], which introduces the use of additional training information in the form of a (labeled) calibration set $(x^{*},\theta^{*})$ that contains observables from the real data distribution as well as the corresponding ground-truth parameters.

### v5

@@ Abstract
Amortized Bayesian inference (ABI) with neural networks can solve probabilistic inverse problems orders of magnitude faster than classical methods. However, ABI is not yet sufficiently robust for widespread and safe application. When performing inference on observations outside the scope of the simulated training data, posterior approximations are likely to become highly biased, which cannot be corrected by additional simulations due to the bad pre-asymptotic behavior of current neural posterior estimators. In this paper, we propose a semi-supervised approach that enables training not only on labeled simulated data generated from the model, but also on unlabeled data originating from any source, including real data. To achieve this, we leverage Bayesian self-consistency properties that can be transformed into strictly proper losses that do not require knowledge of ground-truth parameters. We test our approach on several real-world case studies, including applications to high-dimensional time-series and image data. Our results show that semi-supervised learning with unlabeled data drastically improves the robustness of ABI in the out-of-simulation regime. Notably, inference remains accurate even when evaluated on observations far away from the labeled and unlabeled data seen during training. The code is available at https://github.com/bayesflow-org/self-consistency-real.
@@ 1 Introduction
Theory-driven computational models (mechanistic models) are highly influential across numerous branches of science (Lavin et al., 2021). The utility of computational models largely stems from their ability to fit real data $x$ and extract information about hidden parameters $\theta$. Bayesian methods have been instrumental for this task, providing a principled framework for uncertainty quantification and inference (Gelman et al., 2013). However, gold-standard Bayesian methods, such as Gibbs or Hamiltonian Monte Carlo samplers (Brooks et al., 2011), remain notoriously slow. Moreover, these methods are rarely feasible for fitting complex models (Dax et al., 2021) or even simpler models in big data settings with many thousands of data points in a single dataset (Blei et al., 2017), or when thousands of independent datasets require repeated model re-fits (von Krause et al., 2022).
In recent years, deep learning methods have helped address some of these efficiency challenges (Cranmer et al., 2020). In particular, amortized Bayesian inference (ABI; Gershman and Goodman, 2014; Le et al., 2017; Gonçalves et al., 2020; Radev et al., 2023; Gloeckler et al., 2023; Elsemüller et al., 2024; Zammit-Mangion et al., 2024) has received considerable attention for its potential to automate Bayesian workflows by replacing per-dataset inference with a learned inference model. By training generative neural networks on simulations generated from a probabilistic model, ABI learns a mapping from observations to posterior distributions, allowing subsequent inference on real data to be performed in near-instant time. This amortization trades an upfront computational cost during training for substantial gains in scalability and reusability, making Bayesian inference feasible in settings involving large datasets, intractable likelihood, or real-time constraints.
@@
However, due to the reliance on pre-trained neural networks, ABI methods can become unreliable when applied to data that is unseen or sparsely encountered during training. In particular, posterior samples from amortized methods may deviate significantly from samples obtained with gold-standard MCMC samplers when there is a mismatch between the simulated training data and the real data (Ward et al., 2022; Schmitt et al., 2023; Siahkoohi et al., 2023; Gloeckler et al., 2023; Frazier et al., 2024). This lack of robustness limits the widespread and safe applicability of ABI methods.
@@
In this work, we propose a new robust semi-supervised approach to ABI. The supervised part learns from a “labeled” set of parameters and corresponding synthetic (simulated) observations, $\{\theta,x\}$, while the unsupervised part leverages an “unlabeled” data set of real observations $\{x^{*}\}$ without parameters. In contrast to other methods aiming to enhance the robustness of ABI, our approach does not require ground-truth parameters $\theta^{*}$ (Wehenkel et al., 2024), post hoc corrections (Ward et al., 2022; Siahkoohi et al., 2023), or specific adversarial defenses (Gloeckler et al., 2023), nor does it entail a loss of amortization (Ward et al., 2022; Huang et al., 2023) or generalized Bayesian inference (Gao et al., 2023; Pacchiardi et al., 2024).
While Schmitt et al. (2024) introduced self-consistency (SC) losses to improve simulation efficiency in a fully supervised, simulation-only setting, we extend self-consistency losses to unsupervised training on real observations, enabling robust inference without access to ground-truth parameters. We prove that self-consistency losses are strictly proper, target the true analytic posterior, and are agnostic to the data distribution. These properties allow them to be seamlessly integrated with any simulation-based loss for semi-supervised training without introducing a trade-off in the training objective. Empirical results on diverse tasks, including high-dimensional time series and images, show that our method preserves ABI’s characteristic speed while achieving strong robustness, even with as few as four unlabeled real-world observations. Posterior estimates remain accurate and well-calibrated, generalizing beyond both labeled and unlabeled training data.
@@ 2.1 Bayesian self-consistency
Self-consistency leverages a simple symmetry in Bayes’ rule to enforce more accurate posterior estimation even in regions with sparse data (Schmitt et al., 2024; Ivanova et al., 2024). Crucially, it incorporates the likelihood (when available) or a surrogate likelihood during training, thereby providing the networks with additional information beyond the standard simulation-based loss typically employed in ABI (see below).
Under exact inference, the marginal likelihood is independent of the parameters $\theta$. That is, the Bayesian self-consistency ratio of likelihood-prior product and posterior is constant across any set of parameter values $\theta^{(1)},\dots,\theta^{(L)}$,
@@
However, replacing $p(\theta\mid x)$ with a neural estimator $q(\theta\mid x)$ (and analogously for the likelihood) leads to undesired variance in the marginal likelihood estimates across different parameter values on the right-hand side. Since this variance is a proxy for approximation error, we can directly minimize it via backpropagation along with any other ABI loss to provide further training signal and reduce errors guided by density information. Our proposed semi-supervised approach builds on these advantageous properties.
@@ 2.2 Semi-supervised amortized Bayesian inference
The marginal likelihood identity in Eq. (1) is straightforward, but practically never used in traditional sampling-based methods (e.g., MCMC) because they do not provide a closed-form for the approximate posterior density $q(\theta\mid x)$. In contrast, we can readily evaluate $q(\theta\mid x)$ in ABI when using a neural density estimator that allows efficient density computation (e.g., normalizing flows, Kobyzev et al. (2020)). Thus, we can formulate a family of semi-supervised losses of the form:
@@
where $S$ is a strictly proper score (Gneiting and Raftery, 2007) and $C$ is a self-consistency score. The neural networks to be optimized are a generative model $q$ and (optionally) a summary network $h$ extracting lower dimensional sufficient statistics from the data. We will call the first loss component, $\mathbb{E}_{(\theta,x)\sim p(\theta,x)}\left[S(q(\theta\mid h(x)),\theta)\right]$, the (standard) simulation-based loss, as it forms the basis for standard ABI approaches using simulation-based learning. E.g., this is the maximum likelihood loss for normalizing flows (Kobyzev et al., 2020; Papamakarios et al., 2021) or a vector-field loss for flow matching (Liu et al., 2023; Lipman et al., 2023). We will refer to the second loss component as the (Bayesian) self-consistency loss.
@@
$(q^{*},h^{*})=\operatorname*{argmin}_{q,h}\frac{1}{N}\sum_{n=1}^{N}\left[S(q(\theta_{n}\mid h(x_{n})),\theta_{n})\right]+\lambda\cdot\frac{1}{M}\sum_{m=1}^{M}\left[C\left(\frac{p(x^{*}_{m}\mid\theta_{m})\,p(\theta_{m})}{q(\theta_{m}\mid h(x^{*}_{m}))}\right)\right].$ (3)
where $\lambda$ is the self-consistency weight described in Appendix B. Asymptotically for $N\rightarrow\infty$, that is, for infinite training data generated from the simulator $p(\theta,x)$, a universal density estimator (Draxler et al., 2024) minimizing a strictly proper simulation-based loss (Gneiting and Raftery, 2007) is sufficient to ensure perfect posterior approximation for any data. By this, we mean that the posterior approximation becomes identical to the posterior we would obtain if we could analytically compute $p(\theta\mid x)=p(x\mid\theta)p(\theta)/p(x)$. This analytic posterior is sometimes also referred to as “true” or “correct” posterior (of the specified statistical model). In practice, the posterior is rarely analytic, but we can still verify the accuracy of an approximation by comparing it with the results of a gold-standard approach (if available), such as a sufficiently long, converged MCMC run (Magnusson et al., 2024).
While neural posterior approximation is perfect asymptotically, its pre-asymptotic performance, that is, when training $q(\theta\mid h(x))$ only on a finite amount of simulated data, can become arbitrarily bad (Frazier et al., 2024; Schmitt et al., 2023). For any data $x^{*}$ that is outside the data space implied by $p(\theta,x)$, for instance, when the model is misspecified (see Appendix B for definition), the posterior approximation $q(\theta\mid h(x^{*}))$ may be arbitrarily far away from the analytic posterior $p(\theta\mid x^{*})$. As a result, a simulation-based loss is insufficient to achieve robust ABI in practice. This is where the self-consistency loss comes in: As we will show, the latter greatly improves generalization to atypical data at inference time.
One particular choice for $C$ is the variance (with respect to $\theta\sim p_{C}(\theta)$) of the log Bayesian self-consistency ratio (Schmitt et al., 2024):
@@
which is motivated by theoretical results by Köthe (2023), who show that minimizing the variance of the Bayesian self-consistency ratio also minimizes the KL divergence between true and approximate posterior. $p_{C}(\theta)$ can be any proposal distribution over the parameter space. In practice, two natural choices for $p_{C}(\theta)$ are either the prior $p(\theta)$ or the current approximate posterior $q_{t}(\theta\mid h(x^{*}))$ at training iteration $t$. Because the largest contributions to $C$ are expected to be in high-density regions of the approximate posterior, we choose $q_{t}(\theta\mid h(x^{*}))$ for all of our experiments, see Appendix B for further details on practical considerations.
@@ 2.3 Self-consistency losses are strictly proper
Below, we discuss the strict properness of Bayesian self-consistency losses, showing that they can be used alongside simulation-based losses without altering the target posterior. To simplify the notation, we denote posterior approximators simply as $q(\theta\mid x)$ without considering architectural details such as the use of summary networks $h(x)$. All theoretical results and their proofs remain valid if $x$ is replaced by $h(x)$, provided that the summary network is expressive enough to learn sufficient statistics from $x$.
@@
The proofs of Propositions 1 and 2 are provided in Appendix A. The strict properness extends to semi-supervised losses of the form (2), which combine standard simulation-based and self-consistency losses.
@@
Proposition 3 holds for every instantiation of $C$ that satisfies Proposition 1. The proof of Proposition 3 follows immediately from the fact that the sum of strictly proper losses is strictly proper. Importantly, since Proposition 3 holds independently of $p^{*}(x)$, it holds both in the case of a well-specified model, where $p^{*}(x)=p(x)$, and also in case of any model misspecification or domain shift where $p^{*}(x)\neq p(x)$. That is, there is no trade-off in the semi-supervised loss (2), since both loss components are globally minimized for the same target, that is, the analytic posterior.
In other words, adding the SC loss does not alter the underlying statistical model and continues to target the analytic posterior of that model. As far as we are aware, all other methods aimed at improving the robustness of ABI work by either explicitly or implicitly adjusting the statistical model or the target distribution (moving away from the analytic posterior), thus introducing some form of regularization. In contrast, SC losses should not be interpreted as regularizers: they are strictly proper losses that directly target the analytic posterior of the specified statistical model (i.e., Target 1 in the taxonomy of Elsemüller et al., 2025).
Our proposed semi-supervised loss combines standard simulation-based neural posterior estimation (NPE) and unsupervised self-consistency. Both target the analytic posterior, but in complementary ways: The standard NPE loss facilitates network training on simulated (labeled) data where ground-truth parameters are known. However, it does not ensure posterior accuracy on the empirical data distribution, particularly in regions where simulated data may be sparse or unrepresentative due to model misspecification. The self-consistency loss complements NPE by enforcing Bayes’ rule on observed data and explicitly rewards correct marginal likelihood estimation. This ensures alignment between prior, posterior and likelihood on the empirical data distribution–a property not guaranteed by simulation-based losses alone.
For completeness, we also define strictly proper self-consistency losses for likelihood estimators, rather than posterior approximators.
@@
The proof of Proposition 4 proceeds in the same manner as for Proposition 1, just exchanging likelihood and posterior. Clearly, strict properness does not necessarily hold if both posterior and likelihood are unknown or approximate because any pair of approximators $q(\theta\mid x)$ and $q(x\mid\theta)$ that satisfy $q(\theta\mid x)\propto q(x\mid\theta)\,p(\theta)$ minimize the SC loss. For example, the choices $q(\theta\mid x)=p(\theta)$ and $q(x\mid\theta)\propto 1$ minimize the SC loss, but may be arbitrarily far from their actual target distributions $p(\theta\mid x)$ and $p(x\mid\theta)$, respectively.
In other words, if both likelihood and posterior are unknown, the SC loss has to be coupled with another loss component, such as the maximum likelihood loss, to enable joint learning of both approximators $q(\theta\mid x)$ and $q(x\mid\theta)$. Nevertheless, SC still yields notable improvements: in our experiments, the semi-supervised loss (2) considerably enhanced the robustness of ABI even when both the posterior and likelihood are unknown.
@@ 3 Related work
The robustness of ABI and simulation-based inference methods more generally has been the focus of multiple recent studies (e.g., Frazier et al., 2020; Frazier and Drovandi, 2021; Frazier et al., 2024; Dellaporta et al., 2022; Ward et al., 2022; Cemgil et al., 2020; Gloeckler et al., 2023; Huang et al., 2023; Gao et al., 2023; Siahkoohi et al., 2023; Kelly et al., 2024; Wehenkel et al., 2024; Pacchiardi et al., 2024; Schmitt et al., 2023). These efforts can be broadly classified into two categories: (a) analyzing or detecting simulation gaps and (b) mitigating the impact of simulation gaps on posterior estimates. Since our work falls into the latter category, we briefly discuss methods aimed at increasing the robustness of fully amortized approaches. E.g., Gloeckler et al. (2023) explore efficient regularization techniques that trade off some posterior accuracy to enhance the robustness of posterior estimators against adversarial attacks. Ward et al. (2022) and Siahkoohi et al. (2023) apply post hoc corrections based on real data, utilizing MCMC and the reverse KL-divergence, respectively.
Differently, Gao et al. (2023) propose a departure from standard Bayesian inference by minimizing the expected distance between simulations and observed data, akin to generalized Bayesian inference with scoring rules (Pacchiardi et al., 2024). Perhaps the closest work in spirit to ours is Wehenkel et al. (2024), which introduces the use of additional training information in the form of a (labeled) calibration set $(x^{*},\theta^{*})$ that contains observables from the real data distribution as well as the corresponding ground-truth parameters.